OpenStudy (anonymous):

consider y = √x,find dy/dt,given x=9 and dx/dt=3

OpenStudy (anonymous):

Differentiate implicitly with respect to t. let y = x^(1/2) dy/dt = (1/2)x^(-1/2) dx/dt now substitute for x, y, and dx/dt, and find dy/dt.

OpenStudy (aakashsudhakar):

y = √(x) = x^(1/2) dy/dt = (1/2)(x^(-1/2)) dx/dt = (1/2)(1/√(x)) dx/dt = (1/2√(x)) dx/dt dy/dt = 1/(2√(x)) dx/dt Plug in given values. dy/dt = 1/(2√(9))*(3) dy/dt = 3/(2(±3)) dy/dt = 3/(±6) dy/dt = ±1/2 @Easyaspi314, you mind checking my work on the last bit? I'm a little confused how I could have gotten a ± value for a numerical derivative...

OpenStudy (anonymous):

Correct..except at the end, sqrt(9) = 3 only, not + and - 3. So your answer will change to 1/2 only.

OpenStudy (aakashsudhakar):

Got it, thank you very much! The answer would indeed be dy/dt = 1/2 then.

OpenStudy (anonymous):

Agree.

OpenStudy (anonymous):

thanks all